by TEI TEI

03/05/08 – TTC and Tony figured out the billing issues and TTC sent an email to Grady on how to bill for extra services. The actual project is complete. The County has to fix its own PO to pay Triton but TEI work is complete on this project.

01/25/08 – Tony and I talked today about closeout on this project. Grady is to bill the County the remaining balance as a change order to this job. I will write a change order to match the total amount owed to Triton on this job. The County has a PO invoicing issue on this job (JC bill to this PO).

11/16/07 – TTC talked to Mike F today about project status and sent AKM the following email:

The summer boiler flue temperatures are still no where near 310 F so the fix worked to make things better but we are still probably going to have condensation issues.

The winter boilers are on now and they are slightly better also but we are still not getting 310 F at the outlet of the boiler stack. I told Mike that we should re-smoke test the boiler stack just like we did at the start of the project to see if any drafts are still cooling the boiler stack down as the flue gasses go up.

Mike believes that Triton did a pretty good job buttoning everything on the boiler stack from the job that we bid out and the extra work that you had them do on T&M inside the building. We just aren’t getting as good as results as we would have liked.

The summer boiler is slightly over 2,000,000 btuh which means it would take 4 of the Knight boilers to replace it and meet the same load. We could do this if it was the only problem you still had. The large winter boilers are 5mil btuh so those cannot be replaced by these hot water boilers so we are still kind of faced with some of the same problems with temperatures that you have faced from the beginning in that the outlet temperature at the top of the boiler stack is under 310 F.

and told the County that when they were ready to take the next step to let ThermalTech know and we can help with the design or problem solving.

05/22/07 – Triton submitted their final payapp for the project. There is a $980- change order for the extra camera work. This increases the contract amount and TEI needs approval from the County to pay that. Please advise.

05/14/07 – TEI contacted Triton for closeout documents. Grady said they would send them in ASAP. TEI will return a video copy of the boiler stack video to Grady to closeout the project. TEI has been in contact with MF about the type of brick to replace in the chimney wall to maintain integrity and this would be solid lightweight concrete block according to Spencer at Cole Russell. This project is ready to archive. 12/04/06 – TTC questioned the use of the full allowance today ($5000) and requested an itemized breakdown from Triton for the work completed. The dwgs ask for the the wall to be built by the contractor but that has not been completed to date. The additional work that TEI is aware of is sealing the upper stack and the camera record of the interior. TEI asked for the requested change order to seal the outer stack to be taken from the allowance not as a change order.

11/30/06 – TTC asked about writing a change order to the project for Triton to do the stack sealant work. The project itself is complete. We are moving into the next phase for 2007 where we will address insulation the stack in some way to keep air from overcooling the inner shell. The County is going to open up the wall sections to allow Triton to patch the outer wall of the boiler stack. This can happen right away if the PO gets approved for Tony. TTC asked the County for an update today.

11/13/06 – A lot has happened since the last entry to which i will summarize. Triton has been on site and finished the work. A camera was dropped down between the stacks for a record of how they were built. We discovered that the outer shell had 3 large rectangles removed where spaces were welded to the inner stack. The stack was not support to the outside wall as originally intended. We believe these holes are contributing to the excessive draft up the stack causing a high cooling effect on the boiler flue gas. We instructed Triton to seal the top of the outer stack to the inner stack with a metal seal and put the breeching cover back on seal it air tight with caulk. In the basement Triton sealed the inner stack against the outer stack where the new stack enters the side of the old stack. They did this with steel plates and welds. The temperature sensor is installed and monitoring outlet temperatures and we can only see a 2-degree rise in temperature with the stack being sealed top and bottom. This is not surprising since the are three gaping holes in the outer stack wall on the 4th, 8th and 12th floors that may still be an escape route for air movement. As this is the first stage of the fix it plan being used for investigation as well, we are formulating the second stage now. We want Triton to patch the remaining three holes by change order if necessary. In step two we may try to fill the interstitial space between the stacks with concrete and create a permanent insulation space between the two stacks and this way when the inner stacks rusts out, the concrete stack will remain providing a permanent stack for a long time. We have to get a structural involved at that point to help check the loads on the building.

08/28/06 – We held a construction meeting to get a new handle on where the project stands and to review the R&P chimney stack report. Grady brought one copy of the report and Tony asked for 3 copies to be made and given to the County. The report has basically found that the stack is in FAIR condition. The stack was basically constructed in two different segments, the bottom 130 ft and the top 40 ft. In reviewing the original report from 2001 and the new report from 2006 there is considerable degradation from low temperature condensation and corrosion inside the stack. Photographic evidence was provided in the report to support this claim. The wall thicknesses have decreased and will continue to decrease and the new temperature outlet sensor installed in phase 1 of this project has further provided trending and data to prove that decent outlet temperatures at the boiler flue’s in the basement are significantly reduced in their upward travel to leave the building. Mike Florimonte provided a trend that showed summer boiler leaving temperatures of over 500 F but at the outlet of the stack temperatures were only 210 F. This is a condition that will cause low temperature corrosion. R&P is recommending a spray on coating to protect the interior of the stack. Triton and ThermalTech recommend further investigation into the over cooling effect on the boiler stack. For this much temperature loss to occur, we believe it must be related to the high velocity draft witnessed around the boiler stack at a previous field trip by TE. We have scheduled a Saturday visit with Triton to set off a smoke test inside the boiler stack. We want to witness the path of the released smoke and find the air path that is causing the stack losses. In essence the space between the inner stack and out stack should be dead air space. In addition, Triton will investigate the OA paths surrounding the boiler stack to help ensure that no OA is subjected directly to the boiler stack space. Once this data is resolved, we can isolate the boiler room from the chimney stack and address other “sealant” issues with the new and old breeching. We are scheduled to meet on Sat, Sept 9th. MF will alert the building occupants. AKM requested a synopsis of or plan of attack.

06/16/06 – R&P Industrial Chimney showed up on site. At 2:17 they completed their fieldwork and reported the seems in the stack appear to be the worst off. The took a disposable camera and ultrasonic machine into the stack with them. TTC asked for them to call with a report after they packed up.

05/22/06 – Boiler stack inspection still set for June 19th. Grady is checking on controls and trending.

04/06/06 – Mike called today about getting study complete and temperature sensor in place. Grady replied that the inspection could be the middle of May. The County will need this inspection to be completed either after hours or on the weekend because of hot water reheat. TTC to review and try to find a holiday instead that may work better.

03/09/06 – Pre-construction Mtg was held and notes distributed.

01/17/06 – Wayne reported today that the Bid from Triton did not included a personal property tax statement or a ohio dept statement. TE to check and correct for Wayne.

12/13/05 – Contracts were returned for corrections. TE corrected, printed and returned to the County for submission to the commissioners.

12/01/05 – TTC found the original bids for the job (in sbu’s mailibox) and sent them on to Tony for his use. As soon as contract is approved we can have a pre-construction mtg and get the project moving.

11/08/05 – Tony authorized ThermalTech to write the contract for this work. TE will try to get it done in one day.

10/18/05 – All bids came in and contractors are calling for bid results. TTC gave to Tony and Mike on 10/20 for review. Bid Results – PDF

10/14/05 – No addendum had to be issued. Bids are due next week. TE has received a couple of “No Bids” on this project already.

09/29/05 – Final Project manuals have been created. IES information added for controls. Prebid is set for Oct 6 at 9:30 a.m. TE invited Peck, Triton, Process, Geiler, DK and Langdon to bid the project.

09/09/20/05 – TTC is reviewing Sam’s final drawing and project manual.

08/31/05 – MF reported that with their flow hood the outdoor access door when opened drew 1600 CFM with boiler make-up and exhaust running and 650 cfm with the MUA and EF off. This may contribute to extra cooling down of the boiler flue and TE is investigating welding the outer shell closed during our project. TTC cannot explain the huge draft of air up the shaft at all times

08/30/05 – TTC stopped by the site today to review an access door in the side of the building. There appears to be a huge draft from the outdoors into this access door and it apparently is going up the stack and exiting somewhere that in undetermined. This presents a problem with finalizing our design. We can still go about sealing the basement mechanical room as planned but after we are all done there may still be another step to isolating the inner/outer stacks when we are complete and re-measure the infiltration airflow. In talking with Mike, TE may give Mike the cutout for the future access door. If we cut the whole in the drywall now so we can investigate better how the inner and outer stacks are routed through the building. TTC needs to look at original report and talk to contractor about installation method. There are still unknowns about the original design that are affected an effective solution to this problem.

08/25/05 – TTC sent out final bid manual for review today to MF and AKM. Bid opening might be on 9-7-05.

07/18/05 – TE started the small project manual for bids today. TE has to create a new SOSS because the study used up the first SOSS but since so much of the design is already complete TE only wants to write the new SOSS for $2000 instead of the standard $3500 for bid and contract administration.

06/17/05 – TTC and MF discussed the report today and TTC recommended that we put a small packet together now to close up the air space around the stack. An incremental approach to fixing the problem is best. After the space is enclosed, the stack losses should be re-measured at various conditions to asses the new operating characteristics of the boiler plant. This should be done this year if possible. At the same time, a drywall access hole should be cut to inspect how stacks are anchored together. TE will wait for the County to advise.

06/16/05 – TE sent the final report to Tony and Mike for review and use. TE recommended closing off the bottom of the stack leading into the basement and re-creating a dead air space that can be used for isolation the temperature and creating a thermal barrier.

06/06/05 – TE had a meeting today to discuss recommendations. Our report will be out later this week and 2006 estimate will be estimated based on this report. We plan on discussing with Tony and Mike after Mark Miller sends final report to County. TTC and Mike discussed field findings from last week in an impromptu meeting at B-95 today.

05/24/05 – Mike called today and TTC told him that ThermalTech has not done any construction documents yet on this project and that might be due to the fact that we haven’t given Tony a final report on the condition of the stack (awaiting the re-evaluation from either HC Nutting or the other contractor that MF was going to call into survey the breeching) and TE has not provided a written report to Tony and Mike with recommendations about how to proceed. As MF and TTC discussed the situation though, we decided that if something is to be done this year before heating season, the design and SOSS must begin immediately. TTC will email AKM and ask about how to continue work on this project. TE still has $1172 left from the original SOSS that we do not need to used except to provide a written summary to AKM about our findings.

05/11/05 – TTC gave MF HC Nutting as contact for pipe thickness survey.

03/28/05 – TTC & MM met on site and went over boiler stack calcs and heat losses. Too much heat is being loss going up the stack. We need too look at insulating the stack to keep the heat higher throughout the path.

02/22/05 – TTC and MM reported that initial reports look like stack losses primarily are contributed to the low outlet temperature and condensation issues at the 230 building and not stack sizing or entrainment.

02/10/05 – TE received a signed SOSS and posted it to the website today.

02/01/05 – TTC & Mark Miller visited the site for field investigation and SOSS scope.